OpenStudy (anonymous):

segment lm is midsegment of trapezoid ABCD. AB=48, and DC=88 what is lm?

OpenStudy (anonymous):



OpenStudy (anonymous):

any answer choices?

OpenStudy (anonymous):

@CO_oLBoY

OpenStudy (anonymous):

I NEED STEP BY STEP N IN WORDS

OpenStudy (anonymous):

The midsegment is always the average of the two bases So LM = (AB + DC)/2 LM = (48 +88)/2 LM = 136 / 2 LM = 68